# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

### v4

@@ Abstract
Graph Neural Networks (GNNs) have demonstrated impressive performance across diverse graph-based tasks by leveraging message passing to capture complex node relationships. However, when applied to large-scale real-world graphs, GNNs face two major challenges: First, it becomes increasingly difficult to ensure both scalability and efficiency, as the repeated aggregation of large neighborhoods leads to significant computational overhead; Second, the over-smoothing problem arises, where excessive or deep propagation makes node representations indistinguishable, severely hindering model expressiveness. To tackle these issues, we propose ScaleGNN, a novel framework that adaptively fuses multi-hop node features for both scalable and effective graph learning. First, we construct per-hop pure neighbor matrices that capture only the exclusive structural information at each hop, avoiding the redundancy of conventional aggregation. Then, an enhanced feature fusion strategy significantly balances low-order and high-order information, preserving both local detail and global correlations without incurring excessive complexity. To further reduce redundancy and over-smoothing, we introduce a Local Contribution Score (LCS)-based masking mechanism to filter out less relevant high-order neighbors, ensuring that only the most meaningful information is aggregated. In addition, learnable sparse constraints selectively integrate multi-hop valuable features, emphasizing the most informative high-order neighbors. Extensive experiments on real-world datasets demonstrate that ScaleGNN consistently outperforms state-of-the-art GNNs in both predictive accuracy and computational efficiency, highlighting its practical value for large-scale graph learning.
Index Terms:
Scalable graph neural network, over-smoothing, large-scale graphs.
@@ I Introduction
Graph Neural Networks (GNNs) have become the cornerstone of modern graph representation learning [1, 2], driving advances in applications such as node classification [3], link prediction [4, 5, 6], traffic prediction [7], and recommender systems [8, 9, 10]. The essence of GNNs lies in their iterative message-passing schemes [11, 12], whereby nodes update their representations by aggregating information from their local neighborhoods. This capability enables GNNs to naturally encode node features and capture node correlations.
@@
As digital infrastructures evolve, graphs with millions or even billions of nodes and edges have become commonplace, magnifying the necessity for GNNs that are not only expressive but also computationally scalable [13, 14, 15, 16, 17]. In response, recent research communities have developed a rich array of scalable and deep GNN architectures, with significant innovations focused on balancing model expressiveness, computational efficiency, and training feasibility. One line of research is sampling-based scalable GNN methods such as GraphSAGE [18] and Cluster-GCN [19], which attempt to control computational costs by selecting subsets of neighbors during training, thereby reducing memory overhead and enabling mini-batch processing. These approaches help alleviate the burden of large-scale aggregation but may lose important structural information due to incomplete sampling. Another stream is pre-computation-based scalable approaches (including SGC [20], SIGN [21], and GAMLP [22]), which decouple neighborhood aggregation from parameter learning, allowing feature propagation to be computed in advance and greatly accelerating both training and inference[13, 14]. Meanwhile, a growing body of research has focused on deep and high-order GNNs (e.g., S2GC [23], GBP [24], RpHGNN [25]). They integrate information from multi-hop or high-order neighbors, thereby enriching the receptive field of each node and improving the capture of long-range dependencies. These advances have substantially expanded the applicability of GNNs to web-scale scenarios such as large-scale heterogeneous information networks [15, 16, 17].
@@
Despite these substantial developments, GNNs still face two persistent and intertwined challenges when deployed on large-scale and complex graphs. First, as the number of message-passing layers increases, GNNs suffer from the over-smoothing problem that node representations become overly similar and lose their discriminative power. This issue is particularly severe when incorporating high-order neighbors [26, 27, 28, 29, 30], as their contributions may become redundant or even introduce noise, leading to performance degradation. Although several techniques, such as residual connections [31], skip connections [20, 32], and decoupling propagation [33] from feature transformation, have been proposed to alleviate over-smoothing, they often introduce additional computational complexity or fail to effectively balance local and global information. Second, many traditional GNN architectures struggle with the scalability issue that models face high consumption and computational costs when applied to large-scale graphs [34, 35, 36, 37, 38, 39, 40]. The primary reason is the exponential growth in the number of high-order neighbors, which leads to excessive information aggregation and redundant computations. Some recent works address scalability issues [25] by adopting pre-computation-based techniques, mini-batch training, or sampling strategies. While these approaches improve efficiency, they often sacrifice performance due to incomplete multi-hop neighborhood information. Recently, plug-in modules such as RMask [41] have been introduced to further enhance scalable GNNs. These modules can be seamlessly integrated into frameworks like S2GC and GAMLP, and employ techniques such as noise masking or random walk-based neighbor selection to reduce redundancy and over-smoothing. While plug-in enhancements offer practical improvements, most still fundamentally inherit the aggregation schemes of their base models [26, 27, 28, 29, 30] and often lack adaptive, effective feature fusion of information from different neighborhood hops. Overall, despite advances in sampling-based, pre-computation-based, deep and high-order aggregation, and plug-in modules, scalability and over-smoothing issues still hinder scalable GNNs for large-scale and complex graphs. A unified solution resolving both over-smoothing and scalability remains essential for accurate and efficient large-scale graph learning.
@@
To overcome these challenges, we propose a novel framework towards scalable graph neural networks via adapting high-order neighboring feature fusion named ScaleGNN, which fuses multi-hop graph features to address the over-smoothing issue. Specifically, we introduce a learnable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to adjust the relative importance of different neighborhood orders through trainable weight parameters. By regulating these weights, our approach selectively emphasizes the most informative high-order neighbors while minimizing the influence of less useful ones. Additionally, we introduce a low-order enhanced feature fusion mechanism that adaptively integrates low-order and high-order features based on their relevance to the node classification task. This ensures that our ScaleGNN effectively captures both local and global information, mitigating the over-smoothing problem. Furthermore, we incorporate a selective neighbor masking strategy named Local Contribution Score (LCS) and impose LCS-based sparse constraints to prune less important high-order neighbors, which quantify the relevance of high-order neighbors to target nodes, ensuring that only the most useful neighboring nodes contribute to feature aggregation. To validate the effectiveness of ScaleGNN, we conduct extensive experiments on multiple real-world graph datasets. The results demonstrate that our ScaleGNN consistently outperforms state-of-the-art GNN models in terms of both accuracy and efficiency, providing a scalable and robust solution to scalability and over-smoothing problems in large-scale and complex graph learning.
@@
  • We introduce the selective masking strategy named Local Contribution Score (LCS) to reduce multi-hop redundancy and impose sparse constraints based on LCS to prune unimportant high-order neighbors, significantly reducing computational cost with minimal performance loss.
@@ II-A Graph Neural Network
Figure 1: The overview of the proposed ScaleGNN.
@@ II-B Deep Graph Neural Networks
Despite their success, deep GNNs face several significant challenges. Increasing network depth often leads to the over-smoothing problem, where node representations become indistinguishable and model performance deteriorates. Moreover, as the number of layers grows, scalability and computational complexity become major bottlenecks, especially on large-scale or dense graphs. To mitigate these issues, a variety of techniques such as residual connections, identity mapping, and attention-based neighbor selection have been proposed. In summary, deep GNN research has progressed rapidly, with continuous innovation in model architectures, aggregation mechanisms, and training paradigms. However, challenges such as over-smoothing, scalability, and effective high-order feature utilization still require further study, motivating the design of more adaptive and efficient deep GNN frameworks.
@@ IV-A Adaptive High-order Feature Fusion
To mitigate these issues, we propose a method for constructing distinct adjacency matrices for each neighborhood hop. This allows us to isolate the relationships specific to each hop, ensuring that each neighborhood level contributes uniquely to the node representation. We define the pure $i$-th hop adjacency matrix $\mathbf{A}_{i}$ as the difference between the $(i-1)$-th and $i$-th hops of the adjacency matrix:
@@
These weights are learned through back-propagation, ensuring that the model dynamically adjusts the contribution of each neighborhood hop. The use of the softmax function to optimize these weights ensures that contributions across different hops remain balanced. Finally, we obtain the adaptive high-order fused embedding matrix $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ as follows:
@@
where $d$ is the hidden embedding dimension, $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the input feature matrix, and $\mathbf{W}\in\mathbb{R}^{f\times d}$ is the learnable weight matrix for high-order features. The activation function $\sigma(\cdot)$ introduces non-linearity to the feature transformation.
@@ IV-B Low-order Enhanced Feature Aggregation
where $\beta$ is optimized during training, ensuring the model can adaptively balance the trade-off between local and global information, depending on the specific task and graph structure.
@@ IV-D Joint Optimization
where $\bar{\mathbf{A}}_{i}^{\text{filter}}$ is the filtered high-order adjacency matrix, $\bar{\mathbf{A}}_{i}^{\text{filter}}=\mathbb{I}\left[\mathbf{A}_{i}^{\text{filter}}>0\right]$, where $\mathbb{I}[\cdot]$ denotes the element-wise indicator function, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. This term constrains the model to minimize the number of neighbors retained for each node at each hop.
@@
$\mathcal{L}_{\text{LCS}}=\sum_{v\in\mathcal{V}}\sum_{i=2}^{K}\sum_{u\in\mathcal{N}_{i}^{\text{select}}(v)}\left(1-\text{LCS}(v,u,i)\right)^{2},$ (10)
@@
$\mathcal{L}=\mathcal{L}_{Cla}+\lambda_{1}\mathcal{L}_{\text{LCS}}+\lambda_{2}\mathcal{L}_{SC},$ (11)
@@
7:  $\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i}$
8:  $\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}})$
@@
10:  $\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}^{2}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}})$
@@
23:  $\mathbf{H}=\beta\mathbf{H}_{\text{low}}+(1-\beta)\mathbf{H}_{\text{high}}$
@@ IV-E Time Complexity Analysis
NARS $\mathcal{O}(K|\mathcal{V}|Sf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SeHGNN $\mathcal{O}(K|\mathcal{V}|Pf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
RpHGNN $\mathcal{O}(K|\mathcal{V}|Rf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
@@ V-A2 Baselines
TABLE III: Performance comparison of all models on six datasets. Mi-F1 and Ma-F1 are short for Micro-F1 and Macro-F1. Marker * indicates the results are statistically significant against the best-performed baselines (t-test with p-value $<$ 0.01).
@@
Deep GNN SIGN 72.44 71.97 82.11 80.46 79.53 76.44 71.78 62.99 78.21 70.10 64.32 55.49
S2GC 72.78 72.13 82.45 82.09 79.56 76.67 71.88 63.23 77.09 69.54 64.63 55.90
GBP 72.61 72.08 83.52 82.42 80.56 79.13 71.42 62.41 77.27 69.49 64.66 56.07
GAMLP 72.59 72.06 82.31 80.32 79.11 76.15 71.88 63.15 80.34 71.37 64.53 55.90
GRAND+ 72.48 72.03 82.04 80.23 79.24 76.46 71.43 62.30 77.52 69.70 64.11 55.19
LazyGNN 72.91 72.21 82.45 80.87 79.92 77.03 71.63 62.72 77.60 69.90 64.25 55.38
Scal. GNN NARS 72.41 71.21 81.79 80.30 78.82 75.82 70.75 59.88 78.08 70.20 63.52 54.63
SeHGNN 72.90 72.45 83.04 81.43 79.71 77.70 71.82 62.69 79.68 70.85 64.55 55.86
RpHGNN 73.12 72.54 83.27 82.37 80.60 79.32 72.06 63.34 80.89 71.44 64.95 56.50
Ours ScaleGNNb 72.51 72.74 83.42 82.60 80.58 79.25 71.98 63.11 80.75 71.34 64.97 56.60

### v5

@@ Abstract
Graph Neural Networks (GNNs) have demonstrated impressive performance across diverse graph-based tasks by leveraging message passing to capture complex node relationships. However, when applied to large-scale real-world graphs, GNNs face two major challenges: First, it becomes increasingly difficult to ensure both scalability and efficiency, as the repeated aggregation of large neighborhoods leads to significant computational overhead; Second, the over-smoothing problem arises, where excessive or deep propagation makes node representations indistinguishable, severely hindering model expressiveness. To tackle these issues, we propose ScaleGNN, a novel framework that adaptively fuses multi-hop node features for both scalable and effective graph learning. First, we construct per-hop pure neighbor matrices that capture only the exclusive structural information at each hop, avoiding the redundancy of conventional aggregation. Then, an enhanced feature fusion strategy significantly balances low-order and high-order information, preserving both local detail and global correlations without incurring excessive complexity. To further reduce redundancy and over-smoothing, we introduce a Local Contribution Score (LCS)-based masking mechanism to filter out less relevant high-order neighbors, ensuring that only the most meaningful information is aggregated. In addition, learnable sparse constraints selectively integrate multi-hop valuable features, emphasizing the most informative high-order neighbors. Extensive experiments on real-world datasets demonstrate that ScaleGNN consistently outperforms state-of-the-art GNNs in both predictive accuracy and computational efficiency, highlighting its practical value for large-scale graph learning. The source code is available at https://anonymous.4open.science/r/ScaleGNN.
@@ I Introduction
Graph Neural Networks (GNNs) have become the cornerstone of modern graph representation learning [1, 2, 3, 4, 5], driving advances in applications such as node classification [6], link prediction [7, 8], traffic prediction [9], anomaly detection [10, 11], and recommender systems [12, 13, 14]. The essence of GNNs lies in iterative message-passing schemes [15, 16, 17], whereby nodes update their representations by aggregating information from local neighborhoods [18]. This capability allows GNNs to naturally encode node representations and capture the underlying correlations between nodes.
@@
As digital infrastructures evolve, graphs with millions or even billions of nodes and edges have become commonplace, magnifying the necessity for GNNs that are not only expressive but also computationally scalable [19, 20, 21, 22, 23]. In response, recent research communities have developed a rich array of scalable and deep GNN architectures, with significant innovations focused on balancing model expressiveness, computational efficiency, and training feasibility. One line of research is sampling-based scalable GNN methods such as GraphSAGE [24] and Cluster-GCN [25], which attempt to control computational costs by selecting subsets of neighbors during training, thereby reducing memory overhead and enabling mini-batch processing. These approaches help alleviate the burden of large-scale aggregation but may lose important structural information due to incomplete sampling. Another stream is pre-computation-based scalable approaches (including SGC [26], SIGN [27], and GAMLP [28]), which decouple neighborhood aggregation from parameter learning, allowing feature propagation to be computed in advance and greatly accelerating both training and inference. Meanwhile, a growing body of research has focused on deep and high-order GNNs (e.g., S2GC [29], GBP [30], RpHGNN [31]). They integrate information from high-order neighbors, thereby enriching the receptive field of each node and improving the capture of long-range dependencies. These advances have substantially expanded the applicability of GNNs to web-scale scenarios such as large-scale heterogeneous graphs [19, 20, 21].
@@
Despite these substantial developments, GNNs still face two persistent and intertwined challenges when deployed on large-scale and complex graphs. First, as the number of message-passing layers increases, GNNs suffer from the over-smoothing problem that node representations become overly similar and lose their discriminative power. This issue is particularly severe when incorporating high-order neighbors [32, 33, 34], as their contributions may become redundant or even introduce noise, leading to performance degradation. Although several techniques, such as residual connections [35], skip connections [26, 36], and decoupling propagation [37] from feature transformation, have been proposed to alleviate over-smoothing, they often introduce additional computational complexity or fail to effectively balance local and global information. Second, many traditional GNN architectures struggle with the scalability issue that models face high consumption and computational costs when applied to large-scale graphs [38, 39, 40, 41, 42]. The primary reason is the exponential growth in the number of high-order neighbors, which leads to excessive information aggregation and redundant computations. Some recent works address scalability issues [31] by adopting pre-computation-based techniques, mini-batch training, or sampling strategies. While these approaches improve efficiency, they often sacrifice performance due to incomplete multi-hop neighborhood information. Recently, plug-in modules such as RMask [43] have been introduced to further enhance scalable GNNs. These modules can be seamlessly integrated into frameworks like S2GC and GAMLP, and employ techniques such as noise masking or random walk-based neighbor selection to reduce redundancy and over-smoothing. While plug-in enhancements offer practical improvements, most still fundamentally inherit the aggregation schemes of their base models [32, 33, 34] and often lack adaptive, effective feature fusion of information from different neighborhood hops. Overall, despite advances in sampling-based, pre-computation-based, deep and high-order aggregation, and plug-in modules, scalability and over-smoothing issues still hinder scalable GNNs for large-scale and complex graphs. A unified solution resolving both over-smoothing and scalability remains essential for accurate and efficient large-scale graph learning.
@@
To overcome these challenges, we propose a novel framework towards scalable graph neural networks via adapting high-order neighboring feature fusion named ScaleGNN, which fuses multi-hop graph features to address the over-smoothing issue. Specifically, we introduce a learnable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to adjust the relative importance of different neighborhood orders through trainable weight parameters. By regulating these weights, our approach selectively emphasizes the most informative high-order neighbors while minimizing the influence of less useful ones. Additionally, we introduce a low-order enhanced feature fusion mechanism that adaptively integrates low-order and high-order features based on their relevance to the node classification task. This ensures that ScaleGNN effectively captures both local and global information, mitigating the over-smoothing problem. Furthermore, we incorporate a selective neighbor masking strategy named Local Contribution Score (LCS) and impose LCS-based sparse constraints to prune less important high-order neighbors, which quantify the relevance of high-order neighbors to target nodes, ensuring that only the most useful neighboring nodes contribute to feature aggregation. To validate the effectiveness of ScaleGNN, we conduct extensive experiments on multiple real-world graph datasets. The results demonstrate that ScaleGNN consistently outperforms state-of-the-art (SOTA) GNN models in terms of both accuracy and efficiency, providing a robust solution to scalability and over-smoothing problems in large-scale and complex graphs.
@@
  • We introduce the selective masking strategy named LCS to reduce multi-hop redundancy and impose sparse constraints based on LCS to prune unimportant high-order neighbors, significantly reducing computational cost with minimal performance loss.
@@ II-A Graph Neural Network
Figure 1: The overview of the proposed ScaleGNN. To address the scalability problem, ScaleGNN employs a single convolution process with a hybrid adjacency matrix that fuses features from all different neighbor hops, significantly improving computational efficiency. To mitigate the over-smoothing issue, we introduce a high-order redundant feature masking strategy, selectively filtering out less relevant high-order neighbors, which helps alleviate the over-smoothing issue.
@@ II-B Deep Graph Neural Networks
Despite their success, deep GNNs face several significant challenges. Increasing network depth often leads to the over-smoothing problem, where node representations become indistinguishable and model performance deteriorates. Moreover, as the number of layers grows, scalability and computational complexity become major bottlenecks, especially on large-scale or dense graphs [65]. To mitigate these issues, a variety of techniques such as residual connections, identity mapping, and attention-based neighbor selection have been proposed. In summary, deep GNN research has progressed rapidly, with continuous innovation in model architectures, aggregation mechanisms, and training paradigms. However, challenges such as over-smoothing, scalability, and effective high-order feature utilization still require further study, motivating the design of more adaptive and efficient deep GNN frameworks.
@@ IV-A Adaptive High-order Feature Fusion
To mitigate these issues, we propose a method for constructing distinct adjacency matrices for each neighborhood hop. This allows us to isolate the relationships specific to each hop, ensuring that each neighborhood level contributes uniquely to the node representation. We define the pure $i$-th hop adjacency matrix $\mathbf{A}_{i}$ as the difference between the (i1)(i-1)-th and $i$-th hops of the adjacency matrix:
@@
These weights are learned through back-propagation, ensuring that the model dynamically adjusts the contribution of each neighborhood hop. The use of the softmax function to optimize these weights ensures that contributions across different hops remain balanced. Finally, we obtain the adaptive high-order fused embedding matrix $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d_{f}}$ as follows:
@@
where $d_{f}$ is the hidden embedding dimension, $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the input feature matrix, and $\mathbf{W}\in\mathbb{R}^{f\times d_{f}}$ is the learnable weight matrix for high-order features. The activation function $\sigma(\cdot)$ introduces non-linearity to the feature transformation.
@@ IV-B Low-order Enhanced Feature Aggregation
where $\beta$ is the hyperparameter, ensuring the model can effectively balance the trade-off between local and global information, depending on the specific task and graph structure.
@@ IV-D Joint Optimization
where $\bar{\mathbf{A}}_{i}^{\text{filter}}$ is the filtered high-order adjacency matrix, $\bar{\mathbf{A}}_{i}^{\text{filter}}=\mathbb{I}\left[\mathbf{A}_{i}^{\text{filter}}>0\right]$, where $\mathbb{I}[\cdot]$ denotes the element-wise indicator function, and $\|\cdot\|_{1}$ denotes the $\mathcal{L}_{1}$ normalization. This term constrains the model to minimize the number of neighbors retained for each node at each hop.
@@
$\mathcal{L}_{LCS}=\sum_{v\in\mathcal{V}}\sum_{i=2}^{K}\sum_{u\in\mathcal{N}_{i}^{\text{select}}(v)}\left(1-\text{LCS}(v,u,i)\right)^{2},$ (10)
@@
$\mathcal{L}=\mathcal{L}_{Cla}+\lambda_{1}\mathcal{L}_{LCS}+\lambda_{2}\mathcal{L}_{SC},$ (11)
@@
[…]
@@
[…]
@@
[…]
@@ IV-E Time Complexity Analysis
NARS $\mathcal{O}(K|\mathcal{V}|Sf+K|\mathcal{V}|fd_{f})$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SeHGNN $\mathcal{O}(K|\mathcal{V}|Pf+P^{2}|\mathcal{V}|fd_{f})$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
RpHGNN $\mathcal{O}(K|\mathcal{V}|Rf+R|\mathcal{V}|fd_{f})$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
@@ V-A2 Baselines
TABLE III: Performance comparison of all models on six datasets. Mi-F1 and Ma-F1 are short for Micro-F1 and Macro-F1. Marker * indicates the results are statistically significant against the best-performed baselines (t-test with p-value << 0.01).
@@
Deep GNN SIGN 72.44 71.97 82.11 80.46 79.53 76.44 71.78 62.99 78.21 70.10 62.32 53.59
S2GC 72.78 72.13 82.45 82.09 79.56 76.67 71.88 63.23 77.09 69.54 62.63 53.80
GBP 72.61 72.08 83.42 82.42 80.56 79.13 71.42 62.41 77.27 69.49 62.66 53.97
GAMLP 72.51 72.06 82.31 80.32 79.11 76.15 71.88 63.11 80.34 71.34 62.53 53.79
GRAND+ 72.48 72.03 82.04 80.23 79.24 76.46 71.43 62.30 77.52 69.70 61.85 53.49
LazyGNN 72.91 72.21 82.45 80.87 79.92 77.03 71.63 62.72 77.60 69.90 61.97 53.55
Scal. GNN NARS 72.41 71.21 81.79 80.30 78.82 75.82 70.75 59.88 78.08 70.20 62.52 53.63
SeHGNN 72.90 72.45 83.04 81.43 79.71 77.70 71.82 62.69 79.68 70.85 63.75 54.74
RpHGNN 73.12 72.54 83.27 82.37 80.60 79.32 72.06 63.34 80.89 71.44 63.95 55.50
Ours ScaleGNNb 72.59 72.74 83.42 82.40 80.58 79.15 71.68 62.15 79.75 70.67 63.97 55.60